# Mathematical Modeling and Machine Learning for Predicting Shade-Seeking Behavior in Cows Under Heat Stress

### v1

@@ Abstract.
In this paper we develop a mathematical model combined with machine learning techniques to predict shade-seeking behavior in cows exposed to heat stress. The approach integrates advanced mathematical features, such as time-averaged thermal indices and accumulated heat stress metrics, obtained by mathematical analysis of data from a farm in Titaguas (Valencia, Spain), collected during the summer of 2023. Two predictive models, Random Forests and Neural Networks, are compared for accuracy, robustness, and interpretability. The Random Forest model is highlighted for its balance between precision and explainability, achieving an RMSE of $14.97$. The methodology also employs $5-$fold cross-validation to ensure robustness under real-world conditions. This work not only advances the mathematical modeling of animal behavior but also provides useful insights for mitigating heat stress in livestock through data-driven tools.
@@
Machine Learning Models; Random Forests; Neural Networks; Mathematical Modelling; Heat Stress in Livestock; Precision Livestock Farming
@@ 1. Introduction
Artificial Intelligence (AI) is a field of Computer Science that focuses on creating systems capable of performing tasks that normally require human intelligence. These systems include a wide range of tasks ranging from learning, perception, or reasoning to problem solving, image recognition, or decision making. Moreover, Machine Learning (ML) is an AI discipline that allows a machine, through the use of algorithms, to learn automatically. This enables computers to detect patterns, make predictions, or classify individuals. There is no doubt that AI in general and ML in particular are nowadays applied in practically all areas of knowledge (such as, for instance, in the Economy [5] or Sport Sciences [3]).
In this article, we focus our attention on the application of supervised ML algorithms techniques in the livestock sector. Supervised algorithms are one of the three types of algorithms in the field of ML. The other two are unsupervised learning and reinforcement learning. In the case of supervised learning, we have a (training) dataset including the correct outputs for some inputs. In this way, we intend that the system will automatically be able to obtain the outputs for similar inputs (other than those in the training dataset) with high reliability. This type of database is called a labeled data set. In unsupervised learning, we work with unlabeled dataset. In this case, we try to make the system capable of finding patterns or grouping similar entries. Finally, in reinforcement learning, the model learns to make decisions through trial and error, receiving rewards or penalties in the training process. Of course, in the case of the livestock sector, ML applications already exist to transform and improve animal production and welfare (see, for instance, [1, 25]).
@@
While ML techniques are widely applied in livestock farming, most studies focus on milk production ([10]), disease prediction ([15]), or feed optimization ([1]). However, few efforts have targeted behavioral responses to environmental stress, such as shade-seeking, despite its critical importance for mitigating heat stress in animals. This paper addresses this gap by developing predictive models adapted for livestock management in Mediterranean regions, which are particularly vulnerable to heat stress.
In our case, we will use three well-known algorithms of Supervised Learning: Random Forest (as a generalization of a Decision Tree, which will be the first) and Neural Networks. Random Forest has a range of applications in livestock farming, such as, for instance, in meat quality classification ([24]), fertility and reproduction prediction ([9]) or environmental health management ([27]). Neural Networks have also been widely used in similar contexts (see, for instance, [7, 17, 14]).
In this context, a relevant difference between these two types of algorithms (Random Forest and Neural Networks) lies in the concepts of explainability and interpretability. There are indeed many ML algorithms, for example, in the case of supervised learning, some of the best-known are, Linear Regression, Decision Trees, Logistic Regression, Random Forest, Neural Networks, Gradient Boosting, SVM, KNN, Naive Bayes… Explainability and interpretability are key concepts when it comes to understanding and trusting models. On the one hand explainability refers to the ability of the model to provide understandable explanations of its behaviour and results. In other words, explainability focuses on providing details and reasons about how and why a prediction was produced. On the other hand, interpretability refers to the ease with which a human being can understand the reasons behind a model’s predictions. In this sense, Linear Regression and Decision Trees are interpretable and explainable models (called White Box Models) while Neural Networks need explainability (Black Box Model). In Figure 1 we show a graph with the relationship between the explainability and interpretability of various machine learning supervised ML algorithms. Note that there are techniques (such as LIME or SHA) that can provide explanations for non explainable algorithms. In this article we will not use such techniques. Information about interpretability and explainability can be found in [16].
@@
As mentioned above, in this paper we focus on three models: Decision Trees, Random Forest and Neural Networks. For the sake of completeness, in the following a brief explanation of how these three algorithms work is given. A Decision Tree is a type of supervised learning algorithm used for both classification and regression tasks. It works by splitting a dataset into smaller subsets according to the possible outcomes that may occur depending on decisions. The tree starts at the root node (representing the entire dataset and the first attribute/feature) and splits the data into subsets based on an attribute that maximizes a specific criterion. Each decision node has two or more branches (sub-nodes) based on the possible values of the attribute (see [20, 21]). Decision Trees are very simple to understand and they can be visualized and interpreted easily. However trees can become very complex and overfit the training data and can be biased if some classes dominate (see Section 2 for more information).
To solve this problem, we have the Random Forest algorithms which also works for classification and regression tasks. The Random Forest is an ensemble of multiple Decision Trees combining the predictions of several base estimators to improve robustness and accuracy. It provides estimates of feature importance, which can help in understanding the underlying structure of the data (see [4, 13]). This model is less interpretable compared to a single Decision Tree due to the aggregation of multiple trees and can also be slower to train compared to simpler models, especially with large datasets. More information can also be found in Section 2.
Finally a Neural Network is a model inspired by the human brain’s structure and function. It consists of interconnected layers of nodes (called neurons), where each connection has a weight that adjusts as learning progresses. It can be used to identify patterns and relationships in data through a training process but also for classification and regression. During training, the network learns by adjusting weights based on the errors of its predictions compared to known outcomes (see [6, 12]). As before, more information can be found in Section 2.
@@
We end this first section by talking about the main feature that will determine our mathematical model (and which will be explained at the beginning of the second section): the Temperature-Humidity Index (THI). Mitigating climate change impacts on dairy cattle production systems is essential for the sector’s sustainability, especially as rising temperatures increase the risk of heat stress and negatively impact animal welfare and productivity, particularly in Mediterranean regions. In response, AI integrated with precision livestock farming (PLF) technologies, enables detailed analysis of individual data, supporting the adoption and assessment of targeted strategies to reduce heat stress. The THI index, refers to a measure used to evaluate the combined effects of temperature and humidity on the well-being and performance of animals, particularly livestock such as dairy cattle, poultry, and pigs. It is commonly used in agriculture and animal husbandry to assess the risk of heat stress, which can significantly impact animal health, productivity, and welfare (see [19]). There are different formulas to calculate THI depending on the context, the type of animal and the region (see [8] and the references therein). In our case, the formula we are using, originally proposed by the National Research Council (NRC) in 1971 [18], is:
$\text{THI}=(1.8\times\text{T}_{\text{db}}+32)-(0.55-0.0055\times\text{RH})\times(1.8\times\text{T}_{\text{db}}-26),$ (1.1)
where $\text{T}_{\text{db}}$ is dry bulb temperature in Celsius and RH is the relative humidity in decimal form. To the best of our knowledge, this is the first time that this type of models has been used to predict thermal stress based on behavioral changes in animals and THI data, conducting non-invasive methods for data capture using computer vision.
@@
In this paper we aim to select the most appropriate AI approach to predict the number of cows under the shadow under different THI conditions. To this end, Decision Trees, Random Forest and Neural Networks will be assessed. Animals exposed to high temperatures often exhibit behavioral adaptations to alleviate thermal discomfort. Among these, seeking shade is a primary response, even preferred over other cooling strategies like sprinklers or showers. Of course, providing shaded areas can effectively reduce the thermal load on animals, enhancing their comfort and performance. Moreover, a study by Schütz et al. (2010) highlighted that dairy cattle prioritize access to shade under heat stress conditions, underscoring its importance as a mitigation strategy ([22]).
@@
The article is structured as follows: after this introductory section, we present the methods and results in Section 2. We begin by describing the dataset used in the study, which comes from a farm in Titaguas, Valencia (Spain). The data include observations of the number of cows in the shade, temperature, relative humidity, and time of day, collected during the summer of 2023, from which new variables related to THI were derived. We also explain the use of $5-$fold cross-validation to evaluate model performance.
Next, we describe the three models used, starting with Decision Trees: we explain their hierarchical structure, node-splitting process, and threshold determination. An example Decision Tree with a depth of $3$ is presented, showing the influence of THI, time of day and cumulative THI on cows’ shade-seeking behavior. The optimal tree depth for the dataset and the issue of overfitting are also discussed. The Random Forest model is introduced as an ensemble of multiple Decision Trees, explaining how randomness is added to the tree-building process to improve accuracy and reduce overfitting. The performance of the Random Forest is compared to a single Decision Tree, analyzing error reduction as the number of trees increases. We conclude this section by explaining the architecture of Neural Networks, including layers, neurons, activation functions, and the backpropagation-based training process. We also discuss the importance of optimizing the Neural Network’s architecture, balancing depth, width, and learning rate.
In Section 3, the three machine learning models are compared in terms of accuracy (RMSE), interpretability, and explainability. The pros and cons of each model are analyzed, along with the trade-offs between slightly lower accuracy and higher interpretability. Therefore, we selected the Random Forest model as an intermediate option in terms of RMSE and interpretability to present the conclusions. In particular, the real and predicted values for the number of animals in the shade on several days (corresponding to the first, second and third quartiles in terms of daily RMSE). We show that the Random Forest model captures the general trend of animal movement but misses short-term fluctuations. In the final Section 4, we present the conclusions of the article.
2. Mathematical model and machine learning techniques
2.1. Mathematical model
The mathematical model is derived from a dataset from a farm located in Titaguas, València in Spain. This farm includes a feedlot with a shaded area in the center, monitored by three cameras that count the number of cows within the shaded area. The dataset consists of observations taken every 5-10 minutes during the summer of 2023, spanning from July 11th, 2023 to October 16th, 2023. The observations differ between day and night.
@@ 2.1. Mathematical model
From these data, we have derived new variables for use in the mathematical model, in addition to the number of cows in the shade and the time. Firstly, the time variable has been transformed into a continuous real value for model training. This transformation allows the model to process time as a numerical feature rather than a categorical one. Secondly, with temperature and relative humidity, we calculate the Temperature-Humidity Index (THI) using Equation 1.1. This index serves as a key indicator in our models, quantifying the combined effect of temperature and humidity on the cows’ comfort. Furthermore, we derived two additional variables from the THI: the previous night’s THI, calculated as the average of the previous night, and the accumulated THI, which is the average from the first hour of the day (07:00) up to the current observation time.
Finally, our mathematical model includes the following variables (columns):
@@
In the rest of the section we detail the machine learning techniques used for model development and validation. It presents the performance metrics and insights from the machine learning models applied to predict the number of cows seeking shade in response to different environmental conditions. Each of these models offers different advantages in terms of accuracy and interpretability (see Section 1), and we will analyze their results to determine the most appropriate approach for this problem. More specifically, we begin examining the Decision Tree, a highly interpretable model that provides valuable insights into the key factors that determine cow behavior. We then move on to Random Forests, which combine multiple Decision Trees to improve prediction accuracy and reduce overfitting. Finally, we will discuss the performance of Neural Networks, which offer a more complex and flexible framework for capturing nonlinear relationships between variables but with less interpretability.
2.2. Decision Tree Algorithms
A Decision Tree is based on a hierarchical structure, as shown for example in Figure 3, in which each node can be divided into two (or eventually more) branches following a condition. At the beginning, only the root node exists, where all the (training) dataset belongs to.
Let us explain how the algorithm divides each node. Assume that a particular $N$ node has some observations. The predicted value of the model for that node consists of the mean $\overline{y}_{N}$ of the values of the observations in it. Then, the training squared error can be computed as
@@ 2.2. Decision Tree Algorithms
To achieve this, a variable $j$ and a threshold $t$ must be selected such that the resulting partition creates two new nodes ($N_{1}$ and $N_{2}$) with the lowest possible error, minimizing $E_{N_{1}}+E_{N_{2}}$. This process is repeated for each node that contains more observations than a predefined threshold (in our case 2), or when the node reaches a depth greater than a set value, which represents the number of splits from the root node to the current node. The final nodes, which can no longer be split, are called leaves.
@@
2.3. Random Forest Algorithms
As the performance of a Decision Tree is limited, an ensemble of many of them can be considered to form a Random Forest. This method is particularly effective for improving accuracy while mitigating overfitting. The algorithm works as follows: given an observation $x$, the output of the Random Forest is given by the mean of the prediction of all Decision Trees in the case of regression or the majority vote in classification (see Figure 4).
@@ 2.3. Random Forest Algorithms
2.4. Neural Networks Algorithms
A Fully Connected Neural Network (FCNN) is a model composed by a composition of many layers, as shown in the Figure 6. Each layer consists of a linear transformation and a composition with a nonlinear function. The number of layers, dimensions of each one and the nonlinear functions used on each are fixed as hyperparameters of the model, while the weights of the linear transformation are optimized on the training.
@@ 2.4. Neural Networks Algorithms
Figure 6. Scheme of an example of a FCNN. The input $x$ represents the input data, while $\widehat{y}$ denotes the model’s prediction. The network consists of an input layer, multiple hidden layers, and an output layer.
During training, backpropagation is used to adjust the network weights. This process consists of calculating the model prediction, $\widehat{y}$, for each observation $x$ in the training dataset and comparing it to the true target, $y$. If $\widehat{y}$ does not coincide with $y$, the weights are updated to reduce the prediction error. This is achieved using a gradient-based optimization algorithm that minimizes the squared error $(y-\widehat{y})^{2}$ (SGD or Adam algorithm). Weight updates are propagated backward through the network, layer by layer. Training is repeated for all observations in the training dataset during a fixed number of times, called epochs, or when the error in some test dataset attains a minimum (early stopping parameter). For the best performance of the (Fully Connected) Neural Network model and to avoid variables with larger scales having more influence on the predictions, all variables are scaled so that their mean is 00 and their standard deviation is $1$.
@@
There are several activation functions available for use in Neural Networks ([2], [23]). However, for our FCNN implementation, we have selected the Rectified Linear Unit activation function for hidden layers. This function, mathematically expressed as $\text{ReLU}(x)=\max(0,x)$, allows the network to efficiently model nonlinear relationships in the data. For the output layer, we use a linear activation function defined as $\text{Linear}(x)=x$. This ensures that the output can take any real value, which is suitable for our prediction task.
The primary challenge when working with Neural Networks lies in determining the optimal architecture—balancing depth (number of layers), width (number of neurons per layer), and learning rate—. While deeper and wider networks have the potential to capture more intricate patterns, they also increase the risk of overfitting and may require significantly more computational resources. Hence, selecting the right configuration is critical to achieving both accuracy and efficiency.
@@
3. Results of the model

### v2

# Soft Computing Approaches for Predicting Shade-Seeking Behaviour in Dairy Cattle under Heat Stress: A Comparative Study of Random Forests and Neural Networks

@@ Abstract.
Heat stress is one of the main welfare and productivity problems faced by dairy cattle in Mediterranean climates. In this study, we approach the prediction of the daily shade-seeking count as a non-linear multivariate regression problem and evaluate two soft computing algorithms—Random Forests and Neural Networks—trained on high-resolution behavioral and micro-climatic data collected in a commercial farm in Titaguas (Valencia, Spain) during the $2023$ summer season. The raw dataset ($6907$ daytime observations, $5$-$10$ min resolution) includes the number of cows in the shade, ambient temperature and relative humidity. From these we derive three features: current Temperature–Humidity Index (THI), accumulated daytime THI, and mean night-time THI. To evaluate the models’ performance a $5$-fold cross-validation is also used. Results show that both soft computing models outperform a single Decision Tree baseline. The best Neural Network ($3$ hidden layers, $16$ neurons each, learning rate $=10^{-3}$) reaches an average RMSE of $14.78$, while a Random Forest ($10$ trees, depth $=5$) achieves $14.97$ and offers best interpretability. Daily error distributions reveal a median RMSE of $13.84$ and confirm that predictions deviate less than one hour from observed shade-seeking peaks. These results demonstrate the suitability of soft computing, data-driven approaches embedded in an applied-mathematical feature framework for modeling noisy biological phenomena, demonstrating their value as low-cost, real-time decision-support tools for precision livestock farming under heat-stress conditions.
@@
Mathematical Modelling; Machine Learning; Soft Computing; Random Forests; Neural Networks; Heat Stress in Livestock; Precision Livestock Farming
@@ 1. Introduction
Artificial Intelligence (AI) is a field of Computer Science that focuses on creating systems capable of performing tasks that normally require human intelligence. These systems include a wide range of tasks ranging from learning, perception, or reasoning to problem solving, image recognition, or decision making. Moreover, Machine learning (ML) techniques—supervised, unsupervised and reinforcement—are now standard tools across science and engineering (such as, for instance, in the Economy [5] or Sport Sciences [3], but also in the case of the livestock sector [1, 25]). In this sense, soft computing is the collective term coined by Zadeh (1994) for computational paradigms—fuzzy logic, neural networks, evolutionary algorithms and their hybrids—that trade exactness for robustness and tolerance to uncertainty. Unlike ”hard computing” methods based on exact logic, soft computing approaches are designed to handle noisy, incomplete or vaguely defined data while delivering solutions that are ”good enough” for complex real-world problems ([28]). Biological processes rarely unfold in the orderly, deterministic manner that traditional computational methods assume, but are conditioned by stochastic fluctuations in the environment, sensor noise, and the intrinsic heterogeneity of living organisms. The behavior of dairy cattle is a clear example: even under identical thermal loads individual cows respond differently, and camera-based counts of shade-seeking animals are unavoidably imprecise. Because soft computing was explicitly conceived to ”compute with words, perceptions and uncertain data”, it offers a natural fit for this problem domain. Techniques such as Random Forests and Neural Networks tolerate incomplete or noisy inputs, capture nonlinear interactions without requiring a fully specified physiological model, and produce approximate—but operationally useful—outputs that can drive real-time management decisions. By leveraging this tolerance to imprecision, soft computing models provide robust, low-cost tools for precision livestock farming, where the objective is not perfect prediction of every individual but reliable, adaptive guidance under biologically variable conditions.
@@
While ML techniques are widely applied in livestock farming, most studies focus on milk production ([10]), disease prediction ([15]), or feed optimization ([1]). However, few efforts have targeted behavioral responses to environmental stress, such as shade-seeking, despite its critical importance for mitigating heat stress in animals. This paper addresses this gap by developing predictive models adapted for livestock management in Mediterranean regions, which are particularly vulnerable to heat stress. In our case, we will use three well-known algorithms of supervised learning: Random Forest (as a generalization of a Decision Tree, which will be the first) and Neural Networks. Random Forest has a range of applications in livestock farming, such as, for instance, prediction of milk production ([10]), disease detection ([15]), meat quality classification ([24]), feed optimization ([1]), fertility and reproduction prediction ([9]) or environmental health management ([27]). Neural Networks have also been widely used in similar contexts (see, for instance, [7, 17, 14]). Details of these three soft computing algorithms can be found in Section 2.
@@
In this context, a relevant difference between these two types of algorithms (Random Forest and Neural Networks) lies in the concepts of explainability and interpretability. Explainability and interpretability are key concepts when it comes to understanding and trusting models. On the one hand explainability refers to the ability of the model to provide understandable explanations of its behavior and results. In other words, explainability focuses on providing details and reasons about how and why a prediction was produced. On the other hand, interpretability refers to the ease with which a human being can understand the reasons behind a model’s predictions. Figure 1 positions the three algorithms studied—Decision Tree, Random Forest and Neural Network—along the interpretability-explainability spectrum. Information about interpretability and explainability can be found in [16].
@@
We end this first section by talking about the main feature on which our dataset (that will be explained at the beginning of the second section): the Temperature-Humidity Index (THI). Mitigating climate change impacts on dairy cattle production systems is essential for the sector’s sustainability, especially as rising temperatures increase the risk of heat stress and negatively impact animal welfare and productivity, particularly in Mediterranean regions. In response, AI integrated with precision livestock farming (PLF) technologies enables detailed analysis of individual data, supporting the adoption and assessment of targeted strategies to reduce heat stress. The THI index, refers to a measure used to evaluate the combined effects of temperature and humidity on the well-being and performance of animals, particularly livestock such as dairy cattle, poultry, and pigs. It is commonly used in agriculture and animal husbandry to assess the risk of heat stress, which can significantly impact animal health, productivity, and welfare ([19]). To the best of our knowledge, this is the first time that this type of models has been used to predict thermal stress based on behavioral changes in animals and THI data, conducting non-invasive methods for data capture using computer vision.
@@
In summary, in this paper we therefore aim to select the most appropriate AI approach to predict the number of cows under the shadow under different THI conditions. To this end, Decision Trees, Random Forest and Neural Networks will be assessed. Animals exposed to high temperatures often exhibit behavioral adaptations to alleviate thermal discomfort. Among these, seeking shade is a primary response, even preferred over other cooling strategies like sprinklers or showers. Of course, providing shaded areas can effectively reduce the thermal load on animals, enhancing their comfort and performance. Moreover, a study by Schütz et al. (2010) highlighted that dairy cattle prioritize access to shade under heat stress conditions, underscoring its importance as a mitigation strategy ([22]).
@@
The remainder of the paper describes the dataset and models (Section 2), presents the results (Section 3), discusses their implications (Section 4) and summarises the main conclusions (Section 5).
2. Materials and Methods
2.1. Dataset description and processing
The dataset for our study originates from a farm located in Titaguas, València in Spain. This farm includes a feedlot with a shaded area in the center, monitored by three cameras that count the number of cows within the shaded area. The dataset consists of observations taken every 5-10 minutes during the summer of 2023, spanning from July 11th, 2023 to October 16th, 2023. The observations differ between day and night.
@@ 2.1. Dataset description and processing
From these data, we have derived new variables for use in the models, in addition to the number of cows in the shade and the time. Firstly, the time variable has been transformed into a continuous real value for model training. This transformation allows the model to process time as a numerical feature rather than a categorical one. Secondly, with temperature and relative humidity, we calculate the Temperature-Humidity Index (THI). There are different formulas to calculate THI depending on the context, the type of animal and the region (see [8] and the references therein). In our case, the formula we are using, originally proposed by the National Research Council (NRC) in 1971 ([18]), is:
$\text{THI}=(1.8\times\text{T}_{\text{db}}+32)-\big{(}(0.55-0.0055\times\text{RH})\times(1.8\times\text{T}_{\text{db}}-26)\big{)},$ (2.1)
where $\text{T}_{\text{db}}$ is dry bulb temperature in Celsius and RH is the relative humidity in decimal form. This index serves as a key indicator in our models, quantifying the combined effect of temperature and humidity on the cows’ comfort. Furthermore, we derived two additional variables from the THI: the previous night’s THI, calculated as the average of the previous night, and the accumulated THI, which is the average from the first hour of the day (07:00) up to the current observation time.
The dataset includes the following variables (columns):
@@
In the rest of the section we detail the methods used for model development and validation. It presents the performance metrics and insights from the machine learning models applied to predict the number of cows seeking shade in response to different environmental conditions. Each of these models offers different advantages in terms of accuracy and interpretability (see Section 1), and we will analyze their results to determine the most appropriate approach for this problem. More specifically, we begin examining the Decision Tree, a highly interpretable model that provides valuable insights into the key factors that determine cow behavior. We then move on to Random Forests, which combine multiple Decision Trees to improve prediction accuracy and reduce overfitting. Finally, we will discuss the performance of Neural Networks, which offer a more complex and flexible framework for capturing nonlinear relationships between variables but with less interpretability.
2.2. Soft Computing Decision Tree Algorithm
A Decision Tree is a type of supervised learning algorithm used for both classification and regression tasks. It works by splitting a dataset into smaller subsets according to the possible outcomes that may occur depending on decisions, as shown for example in Figure 3. The tree starts at the root node (representing the entire dataset and the first attribute/feature) and splits the data into subsets based on an attribute that maximizes a specific criterion.
Let us explain how the algorithm divides each node ([20, 21]). Assume that a particular $N$ node has some observations. The predicted value of the model for that node consists of the mean $\overline{y}_{N}$ of the values of the observations in it. Then, the training squared error can be computed as
@@ 2.2. Soft Computing Decision Tree Algorithm
To achieve this, a variable $j$ and a threshold $t$ must be selected such that the resulting partition creates two new nodes ($N_{1}$ and $N_{2}$) with the lowest possible error, minimizing $E_{N_{1}}+E_{N_{2}}$. This process is repeated for each node that contains more observations than a predefined threshold (in our case 2), or when the node reaches a depth greater than a set value, which represents the number of splits from the root node to the current one. The final nodes, which can no longer be split, are called leaves.
@@
2.3. Soft Computing Random Forest Algorithm
As the performance of a Decision Tree is limited, an ensemble of many of them can be considered to form a Random Forest—which also works for classification and regression tasks—. This method is particularly effective for improving accuracy while mitigating overfitting. The Random Forest is an ensemble of multiple Decision Trees combining the predictions of several base estimators to improve robustness and accuracy. It provides estimates of feature importance, which can help in understanding the underlying structure of the data (see [4, 13]). This model is less interpretable compared to a single Decision Tree due to the aggregation of multiple trees and can also be slower to train compared to simpler models, especially with large datasets.
The algorithm works as follows: given an observation $x$, the output of the Random Forest is given by the mean of the prediction of all Decision Trees in the case of regression or the majority vote in classification (see Figure 4).
@@ 2.3. Soft Computing Random Forest Algorithm
2.4. Soft Computing Neural Networks Algorithm
A Neural Network is a model inspired by the human brain’s structure and function. It consists of interconnected layers of nodes (called neurons), as shown in the Figure 6, where each connection has a weight that adjusts as learning progresses. It can be used to identify patterns and relationships in data through a training process but also for classification and regression. During training, the network learns by adjusting weights based on the errors of its predictions compared to known outcomes (see [6, 12]). Each layer consists of a linear transformation and a composition with a nonlinear function. The number of layers, dimensions of each one and the nonlinear functions used on each are fixed as hyperparameters of the model, while the weights of the linear transformation are optimized on the training.
@@ 2.4. Soft Computing Neural Networks Algorithm
Figure 6. Scheme of an example of a (Fully Connected) Neural Network. The input $x$ represents the input data, while $\widehat{y}$ denotes the model’s prediction. The network consists of an input layer, multiple hidden layers, and an output layer.
During training, backpropagation is used to adjust the network weights. This process consists of calculating the model prediction, $\widehat{y}$, for each observation $x$ in the training dataset and comparing it to the true target, $y$. If $\widehat{y}$ does not coincide with $y$, the weights are updated to reduce the prediction error. This is achieved using a gradient-based optimization algorithm that minimizes the squared error $(y-\widehat{y})^{2}$ (SGD or Adam algorithm). Weight updates are propagated backward through the network, layer by layer. Training is repeated for all observations in the training dataset during a fixed number of times, called epochs, or when the error in some test dataset attains a minimum (early stopping parameter). For the best performance of the (Fully Connected) Neural Network (FCNN) model and to avoid variables with larger scales having more influence on the predictions, all variables are scaled so that their mean is 00 and their standard deviation is $1$.
@@
There are several activation functions available for use in Neural Networks ([2, 23]). However, for our FCNN implementation, we have selected the Rectified Linear Unit activation function for hidden layers. This function, mathematically expressed as $\text{ReLU}(x)=\max(0,x)$, allows the network to efficiently model nonlinear relationships in the data. For the output layer, we use a linear activation function defined as $\text{Linear}(x)=x$. This ensures that the output can take any real value, which is suitable for our prediction task.
The primary challenge when working with Neural Networks lies in determining the optimal architecture—balancing depth (number of layers), width (number of neurons per layer), and learning rate—. While deeper and wider networks have the potential to capture more intricate patterns, they also increase the risk of overfitting and may require significantly more computational resources. Hence, selecting the right configuration is critical to achieving both accuracy and efficiency. To rigorously identify the architecture, 45 different neural networks have been tested with all the combination of these hyperparameters: 1, 3 or 5 layers with 4, 16, 64, 256 or 1024 neurons each and a learning rate of $10^{-2}$, $10^{-3}$ or $10^{-4}$.
@@
3. Results
3.1. Global model performance